1. Liv Enterprises is considering a national launch of 2 corn chip products under project A and B. The national launch will require the following investments:

Additional manufacturing equipment; Project A: K900,000.00 and Project B: K1,000,000.00

Upgrading existing facilities; Project A: K100,000.00 and project B: K400,000.00 Both of the above would be paid for at the outset and would be depreciated in the accounts over a five year period using straight line with a residual value for both Project A and B of zero.

Projected revenues and costs over a period of five years are given in table below:

                                                                        PROJECT A

YEAR

REVENUE

COSTS

1

K1200000

1340000

2

K2000000

K1670000

3

K2000000

K1520000

4

K2250000

K1685000

5

K2250000

K1685000

                                                                         PROJECT B

YEAR

REVENUE

COSTS

1

K1300000

K1460000

2

K2100000

K1520000

3

K2200000

K1400000

4

K2400000

K160000

5

K2500000

K1720000

Required:

(i) Appraise the two projects through the use of Accounting Rate of Return and suggest which of the two offers a better option. Please note that the above costs do not include depreciation.

(ii) results increasingly suggest that Discounted Cash Flow (DCF) techniques are a better way to appraise investment projects as compared to payback method. Summarize reasons why this might be so.

2. In the procurement agreements so signed in question 1 above, explain and give four legal reasons through which these 2contracts can be discharged or terminated.
Answer #1

Live enterprises

(i) Accounting rate of return = (Average net annual income / initial investment) * 100

particulars Project A

Project B

Revenue (average of all 5 years / 5 ) = A (1200000 + 2000000+ 2000000 + 2250000+2250000 ) / 5 = 1940000 (1300000 +2100000 + 2200000 + 2400000 + 2500000 )/5 = 2100000
Cost (average of all 5 years / 5 ) = B (1340000+1670000+1520000+1685000+1685000)/5= 1580000 (1460000+1520000+1400000+160000+1720000)5=1252000
Annual savings = (A-B) = C 360000 848000
Less : depreciaton = (Additional manufacturing equipment) / 5 years = D 900000/5=180000 1000000/5=200000
tax - -
Annual savings (C-D) = E 180000 648000
Add : depreciation 180000 200000
Annual cash inflows 360000 848000

Accounting rate of return = (Average net annual income / initial investment) * 100

Accounting rate of return = (E / initial investment) * 100

  • Project A : (180000/ 900000) * 100 = 20%
  • Project B : (648000/1000000) * 100 = 64.8%

Decision : project B is better .

[note : ARR can be computed alternatively by taking initial investment as a basis of computation { Accounting rate of return = (Average net annual income / average investment) * 100 }. The value of project A and project B would then change accordingly as 40% and 129.6% ]

(ii) Discounted Cash Flow (DCF) techniques are a better way to appraise investment projects as compared to payback method.Reasons are listed below -

  • DCF can be used to calculate the internal rate of return Internal Rate Of Return (IRR) of an investment
  • The discounted payback period accounts for the "time value of money" whereas in case of payback period, it completely ignores it ignores the time value of money

2. Causes for Terminating / discharging Legal Contracts -

  • ​​​​​​​Impossibility of Performance- A contract typically requires one or more parties to do something, which is called performance. . If for some reason it has become impossible , it is called impossibility of performance. The company has the right to terminate the contract in the case of an impossibility of performance.
  • Rescission of the Contract - A rescission of a contract is when a contract is terminated because an individual misrepresented themselves, acted illegally or made a mistake.

  • Breach of Contract - When a contract is intentionally not done by one party, it is breach of contract and is grounds for contract termination. A breach of contract may exist because one party failed to meet his obligations at all or did not meet his obligations fully.

  • Prior Agreement - You may terminate a contract if you and the other party have a prior written agreement that calls for a contract termination because of a specific reason

  • Performance - all the terms of the contract must be precisely completed to discharge liability.